# Pivoting technique for the circle homeomorphism group

### v1

@@ 5. Pivoting technique
When $S^{\operatorname{\mathbb{Z}}_{>0}}\times S^{\operatorname{\mathbb{Z}}_{>0}}\times S^{\operatorname{\mathbb{Z}}_{>0}}$ is endowed with the product measure of $\mu$, #$\#P_{n}$ dominates $X_{1}+\ldots+X_{n}$ in distribution for each $n$. That means,
(#$\operatorname{\mathbb{P}}(\#P_{n}(s)\geq T)\geq\operatorname{\mathbb{P}}(X_{1}+\ldots+X_{n}\geq T)\quad(\forall T\in\mathbb{Z}_{\geq 0}).$
@@
(5.5) (#$\operatorname{\mathbb{P}}\left(\#P_{k+1}(s)\geq i+j\,\Big{|}\,\#P_{k}(s)=i\right)\geq\left\{\begin{array}[]{cc}1-\frac{4\zeta^{2}}{N}&\textrm{if}\,\,j=1,\\ 1-\left(\frac{4\zeta^{2}}{N}\right)^{-j+1}&\textrm{if}\,\,j<0.\end{array}\right.$
Let us prove that for each $k=1,\ldots,n$ and for each $i\in\operatorname{\mathbb{Z}}_{\geq 0}$ we have (#$\operatorname{\mathbb{P}}(\#P_{k}\geq i)\geq\operatorname{\mathbb{P}}(X_{1}+\ldots+X_{k}\geq i)$. For $k=1$, the claim follows from Inequality 5.5 because #$\#P_{k-1}=0$ always. Now, assuming the statement for $k$ as an induction hypothesis, we observe
(#$\operatorname{\mathbb{P}}(\#P_{k+1}\geq i)$ (#$\geq\operatorname{\mathbb{P}}(\#P_{k}+X_{k+1}\geq i)=\sum_{j}\operatorname{\mathbb{P}}(\#P_{k}\geq j)\operatorname{\mathbb{P}}(X_{k+1}=i-j)$
@@
(#$\operatorname{\mathbb{P}}\big{(}\#P_{n}(\mathbf{r},\mathbf{s},\mathbf{t};\mathbf{w})\leq n/2\big{)}\leq\big{(}3\sqrt[4]{4\zeta^{2}/N}\big{)}^{n}\leq 0.6^{n}$
@@
$\big{[}(\mathbf{r},\mathbf{s},\mathbf{t})\sim_{n}^{\prime}(\bar{\mathbf{r}},\bar{\mathbf{s}},\bar{\mathbf{t}})\big{]}\Leftrightarrow\left[\begin{array}[]{c}\textrm{$r_{i}=\bar{r}_{i}$ and $t_{i}=\bar{t}_{i}$ for each $i\in\operatorname{\mathbb{Z}}_{>0}$},\\ \textrm{ $\bar{s}_{i}=s_{i}$ unless $i$ is among the $n/2$ smallest pivotal times of $P_{n}(\mathbf{r},\mathbf{s},\mathbf{t};\mathbf{w})$}\end{array}\right]$
@@
$w_{0}^{\prime}$ $:=W_{i(1)-1}r_{i(1)}=w_{0}\cdot r_{1}s_{1}t_{1}w_{1}\cdots r_{i(1)-1}s_{i(1)-1}t_{i(1)-1}w_{i(1)}r_{i(1)},$
$w_{l}^{\prime}$ $:=t_{i(l)}w_{i(l)}W_{i(l)}^{-1}W_{i(l+1)-1}r_{i(l+1)}$
@@
$w_{\lfloor n/2\rfloor}^{\prime}$ $:=t_{i(\lfloor n/2\rfloor)}w_{i(\lfloor n/2\rfloor)}W_{i(\lfloor n/2\rfloor)}^{-1}W_{n}$

### v2

@@ 5. Pivoting technique
When $S^{\operatorname{\mathbb{Z}}_{>0}}\times S^{\operatorname{\mathbb{Z}}_{>0}}\times S^{\operatorname{\mathbb{Z}}_{>0}}$ is endowed with the product measure of $\mu$, $P_{n}$ dominates $X_{1}+\ldots+X_{n}$ in distribution for each $n$. That means,
$\operatorname{\mathbb{P}}(\#P_{n}(s)\geq T)\geq\operatorname{\mathbb{P}}(X_{1}+\ldots+X_{n}\geq T)\quad(\forall T\in\mathbb{Z}_{\geq 0}).$
@@
(5.5) $\operatorname{\mathbb{P}}\left(\#P_{k+1}(s)\geq i+j\,\Big{|}\,\#P_{k}(s)=i\right)\geq\left\{\begin{array}[]{cc}1-\frac{4\zeta^{2}}{N}&\textrm{if}\,\,j=1,\\ 1-\left(\frac{4\zeta^{2}}{N}\right)^{-j+1}&\textrm{if}\,\,j<0.\end{array}\right$.
Let us prove that for each $k=1,\ldots,n$ and for each $i\in\operatorname{\mathbb{Z}}_{\geq 0}$ we have $\operatorname{\mathbb{P}}(\#P_{k}\geq i)\geq\operatorname{\mathbb{P}}(X_{1}+\ldots+X_{k}\geq i)$. For $k=1$, the claim follows from Inequality 5.5 because $P_{k-1}=0$ always. Now, assuming the statement for $k$ as an induction hypothesis, we observe
$\operatorname{\mathbb{P}}(\#P_{k+1}\geq i)$ $\geq\operatorname{\mathbb{P}}(\#P_{k}+X_{k+1}\geq i)=\sum_{j}\operatorname{\mathbb{P}}(\#P_{k}\geq j)\operatorname{\mathbb{P}}(X_{k+1}=i-j)$
@@
$\operatorname{\mathbb{P}}\big{(}\#P_{n}(\mathbf{r},\mathbf{s},\mathbf{t};\mathbf{w})\leq n/2\big{)}\leq\big{(}3\sqrt[4]{4\zeta^{2}/N}\big{)}^{n}\leq 0.6^{n}$
@@
[…]
@@
$w_{0}^{\prime}$ $=W_{i(1)-1}r_{i(1)}=w_{0}\cdot r_{1}s_{1}t_{1}w_{1}\cdots r_{i(1)-1}s_{i(1)-1}t_{i(1)-1}w_{i(1)}r_{i(1)},$
$w_{l}^{\prime}$ $=t_{i(l)}w_{i(l)}W_{i(l)}^{-1}W_{i(l+1)-1}r_{i(l+1)}$
@@
$w_{\lfloor n/2\rfloor}^{\prime}$ $=t_{i(\lfloor n/2\rfloor)}w_{i(\lfloor n/2\rfloor)}W_{i(\lfloor n/2\rfloor)}^{-1}W_{n}$